OpenStudy (anonymous):

Y varies inversely with x. If y = 24 when x = 48, find x when y = 70. Round to the nearest tenth.

OpenStudy (anonymous):

Please someone answer my question.

OpenStudy (anonymous):

16.5

OpenStudy (anonymous):

y = k/x 24 = k / 48 k = 1152 70 = 1152 / x x = 16.457